OpenStudy (anonymous):

Having issues with this one: Divide: 2x^2-8x/4x-9 / x-4/4x-9

OpenStudy (anonymous):

I have the first part done but then got lost.

OpenStudy (anonymous):

You can cancel out the 4x-9s so you get 2x^2-8x/x-4

OpenStudy (anonymous):

Uhoh where did the /x-4 come from?

OpenStudy (anonymous):

isn't that what you wrote 2x^2-8x/4x-9 / **x-4**/4x-9

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

\[\frac{2x ^{2}-8x }{ 4x-9 }\div \frac{ x-4 }{ 4x-9 }=\frac{ 2x ^{2}-8x }{ 4x-9 }\times \frac{ 4x-9 }{ x-4 }\]

OpenStudy (anonymous):

the 4x-9's go away

OpenStudy (anonymous):

I understand that part

OpenStudy (anonymous):

cancel each other out

OpenStudy (anonymous):

Then you just divide or do you not know the dividing part?

OpenStudy (anonymous):

2x^2-8x/x-4

OpenStudy (anonymous):

do you know how to divide that?

OpenStudy (anonymous):

2x^2/x-4 -8x/x-4 x-4 cancel out I dont know

OpenStudy (anonymous):

you use long division. I will show you just a sec

OpenStudy (anonymous):

k

OpenStudy (anonymous):

It's basically the same as regular division.